OpenStudy (anonymous):

**Please help!! I give medals!! How many solutions does a linear system have, if the lines are intersecting? 0 1 2 infinitely many

OpenStudy (ybarrap):

There is a single intersection, there is a single solution.

OpenStudy (ybarrap):

That is what a solution means.

OpenStudy (nurali):

One, The point where the lines meet is the answer.

OpenStudy (ybarrap):

If I have two lines y = (1/2) x + 3 and and y = (1/3) x - 35, these lines will intersect. I know this because the slopes are different. I know when two lines DON'T intersect, their slopes are the same. Contrapositively, if the slopes are different then the lines DO intersect.

random231 (random231):

@Hannnnah didnt u follow my answer in the previous question?

OpenStudy (anonymous):

Thanks guys and sorry @random231 Im not even reading them Im trying to do this fast because I have work